OpenStudy (anonymous):

Please ! check my answer. A certain game has two stacks of 30 tiles with pictures on them. In the first stack of tiles, there are 10 dogs, 4 cats, 5 balls, and 11 horses. In the second stack of tiles, there are 3 flowers, 8 fish, 12 balls, 2 cats, and 5 horses. The top tile in each stack is chosen. Find each probability.

OpenStudy (anonymous):

1) P (each a ball): First stack 5 balls, second 12 balls (5/30)*(12/30) 2) P (exactly one is a ball): 5/30=1/6

OpenStudy (phi):

I interpret the second question to be what is the probability we get only 1 ball after taking a tile from each stack. The probability we get 1 ball from the first stack and 0 balls from the 2nd stack is (5/30)*(18/30)= (1/6)(3/5)= 1/10 Also, the probability of getting 0 balls from the first stack, and 1 from the 2nd is (5/6)*(12/30)=1/3 Add to get the probability of 1 ball= 13/30 Another way to do it is to note that the probability of 1 ball is 1 - P(2 balls) - P(0 balls) I leave it to you to see you get the same answer.

OpenStudy (anonymous):

1) second stack has 12 balls, how you get 18?

OpenStudy (phi):

The first calculation is P(ball 1st stack)*P(not a ball from 2nd stack)

OpenStudy (phi):

Hopefully 30-12= 18